# Generative AI and Firm Productivity:

### v1

@@ 3.4 Data and Estimation
For the experiments operated at the consumer level, we recorded each consumer’s treatment status. We gathered consumers’ complete set of activities, including the number of product views (referred to as View), product clicks (Click), product orders (Order), and total expenditures on those orders (Sales). [^13] In the Search Query Refinement experiment, product views represent the number of products a consumer browses on the search results page, which displays a summarized collection of products immediately after a query search. In the Google Advertising Title experiment, product views refer to the number of views of advertised products within the Google Shopping tab. In the Search Query Refinement and Product Descriptions experiments, product clicks capture the number of times consumers clicked into product detail pages. In the Marketing Push Message experiment, product clicks reflect consumer clicks on push notifications, while in the Google Advertising Title experiment, they indicate clicks on advertised products. For comparability across workflows, our benchmark analysis focuses on two primary outcome measures: (i) total consumer expenditure (Sales), which serves as a revenue-based measure of retail productivity; and (ii) the conversion rate, a binary indicator for whether a consumer made a purchase, representing a proxy for changes in consumer experience. Both sales and conversion rate are widely used industry metrics. For the product-level experiment, we collected analogous data at the product level. Table 3 presents summary statistics for the key variables in each of the five experiments where granular data are available.
@@
Conversion Rate 0.04 0.20 0 0 1
Sales 0.51 4.56 0 0 2,941
@@
Click 0.058 0.234 0 0 1
Conversion Rate 0.0014 0.037 0 0 1
Order 0.0015 0.039 0 0 6
Sales 0.045 1.756 0 0 501
@@ 6 Chargeback Defense 15% defense success rate increase
The remaining four workflows with detailed data reported in Table 4 exhibit smaller effects on sales, ranging from a negative and statistically insignificant change to gains of up to 3%. Specifically, the Search Query Refinement application increased sales by 2.93% ($p<0.05$), while the Product Description generated a 2.05% ($p<0.05$) gain—still substantial effects for a platform of this scale and maturity. The Marketing Push Message workflow shows a positive yet not statistically significant improvement in sales (1.6%). This likely reflects the combination of a very low baseline conversion rate (only 0.14% of consumers make a purchase) and high variance in expenditures among converters, suggesting that broader implementation could provide sufficient power to detect treatment effects. [^15] Note that this case exhibits significant conversion rate increase in Table 4 as well as significant order increase in Appendix Table C5. It is also important to note that only a subset of the treatment group was exposed to GenAI-generated messages in this experiment, which may attenuate the observed treatment effects under partial exposure. By contrast, the Google Advertising Title workflow exhibits an insignificant negative effect. This pattern is consistent with the lack of fine-tuning of the GenAI model to the advertising context, which led it to omit commonly used commercial keywords. An alternative explanation is a divergent delivery bias in ad distribution: Google’s advertising algorithm may have recognized and deprioritized AI-generated titles, reducing their visibility. Together, these findings underscore the importance of domain-specific fine-tuning or retraining of foundation models when applying GenAI to industry-specific tasks requiring specialized knowledge [27].
@@
These results provide new evidence on the potential of GenAI to enhance revenue-based productivity in online retail, thereby contributing to the broader debate on the economic consequences of GenAI adoption [56, 1]. A key distinction of our study is its focus on revenue outcomes rather than input-side efficiency gains, which makes our estimates not directly comparable to those from studies emphasizing labor productivity or task efficiency. Among the seven deployments we examine, five deliver measurable performance gains, showing that GenAI can generate substantial improvements in firm outcomes under real-world operating conditions. Overall, the evidence points to a positive but heterogeneous impact of GenAI along the customer journey.
@@ 4.2 Mechanism
Most of the existing literature attributes the productivity-enhancing potential of GenAI to supply-side mechanisms, such as labor savings or efficiency improvements (e.g., reductions in the time required to complete a task). Our context and findings highlight an additional channel: GenAI can unlock productivity gains by enriching the demand-side consumer experience. We measure this mechanism using conversion rates (i.e., the likelihood that consumers complete a purchase), which serves as a widely recognized industry standard for consumer satisfaction in the e-commerce context. Across the various workflows, we document significant increases in conversion rates ranging from 1% to 22% (Column 4 of Table 4), which in turn translate into higher output as measured by sales. Notably, we do not find any significant effects along the intensive margin. Table 5 shows that the average cart value among consumers who made at least one purchase (or among products purchased at least once) remains unchanged following GenAI adoption. In our setting, GenAI primarily drove market expansion by converting a larger share of consumers, rather than inducing existing buyers to purchase higher-priced or larger quantities of products.
@@
    The table presents the treatment effects of GenAI adoption on “Cart Value”. For workflows 1-4, “Cart Value” refers to the expenditure per consumer, conditional on the consumer making a purchase. For workflow 5, “Cart Value” refers to the expenditure per product, conditional on the product being purchased.
@@
    Columns (1) report the estimated coefficient, with standard errors in brackets. Columns (2) report % Change, which is calculated dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
This evidence indicates that the observed productivity gains arise from GenAI’s capacity to reduce market frictions—both by enabling new services and by improving existing ones—thereby enhancing the shopping experience and influencing purchasing behavior. For example, GenAI mitigates information asymmetries by providing relevant and timely assistance through pre-sale chatbots (a 21.7% increase in conversion rate, $p<0.01$) and by generating comprehensive and structured product descriptions (a 1.27% increase, $p<0.05$). It reduces search frictions and improves match quality by enhancing the translation and semantic comprehension of consumer queries (a 1.15% increase, $p<0.05$). In addition, GenAI enables personalization of marketing content by enabling large-scale generation of customized marketing messages across a broad product portfolio (a 3% increase, $p<0.05$). Finally, evidence from Chargeback Defense and Live Chat Translation points to improvements consistent with enhanced user experience, including higher success rates and greater consumer satisfaction.
@@ 4.3 Aggregate Productivity Gains Across Workflows
1.6
1.3
2 Search Query Refinement 0.0648 40.6 2.6
3 Product Description 0.0104 52.1 0.5
4 Marketing Push Message 0.0004 365.0 0.1
Total (linear additivity) 5.0 (upper-bound)
@@
  • Column (1) reports the absolute lifts in sales (treatment effects of GenAI) for each workflow from Table 4. Column (2) shows the factor used to annualize the workflow-specific estimates. Column (3) reports the annualized values.
@@
Table 6 summarizes the key variables used in our calculations. Column (1) reports the AI-driven incremental value per consumer for each experimental workflow, reflecting the estimated average treatment effects (the absolute sales lift per consumer) reported in Column (1) of Table 4. Because each experiment had a different duration, Column (2) presents the time multiplier used to extrapolate these effects to an annual horizon. For instance, the Pre-Sale Service Chatbot experiment spans two months, yielding a time multiplier of six (i.e., $12/2$). For each worklow, we obtain the annualized AI-driven incremental value per consumer in Column (3) by multiplying the estimated effect in Column (1) by the corresponding time multiplier in Column (2). This calculation assumes that the treatment effects observed during the experimental period remain constant over time, abstracting from potential amplification (e.g., greater engagement and purchases as GenAI-generated content enhances consumer satisfaction and platform loyalty) or attenuation (e.g., consumer dissatisfaction and product returns due to potential mismatches between AI-generated content and actual product characteristics).
@@
Finally, we aggregate the annualized estimates across the four workflows to obtain the total annual incremental value per consumer attributable to GenAI (see “Total” in Column (3)). This aggregation assumes that effects across workflows are linearly additive and abstracts from potential cross workflow interactions, such as cannibalization among touchpoints or synergies. Based on the four GenAI applications with positive sales effects, we estimate an annual incremental value of about \$5 per consumer, which decreases to \$4.6 when applying the lower-bound estimate from the Pre-Sale Service Chatbot experiment. These effects represent roughly 5.5-6% of the increase in revenue per user observed in global e-commerce between 2023 and 2024 [59], highlighting the economic significance of these gains relative to broader industry trends.
@@
It is worth noting that these estimates capture only a partial and time-dependent view of the firm’s efforts to scale up GenAI across workflows. While in 2023 the platform applied GenAI to only a handful of workflows, by 2024 it was deployed in more than 40 applications and by 2025 in over 60. This rapid expansion is also reflected in the growth of API calls to large language models: in mid-2024, AI-related API requests averaged over 50 million per day, rising to more than 1 billion per day by mid-2025—a twentyfold increase. Hence, our point estimates should be interpreted with caution and in light of these rapid adoption trends, which nonetheless indicate that the firm anticipates substantial value from GenAI.
@@ 5 Heterogeneous Treatment Effects
To further shed light on the mechanism behind our results, we examine heterogeneity in treatment effects across sellers, buyers, and products. If GenAI primarily reduces frictions on both the demand and supply sides, one would expect relatively larger gains among participants with lower baseline capabilities—namely, smaller and less experienced sellers, buyers with limited platform engagement, and products in the long tail of sales or in less concentrated categories. We classify each dimension into “high” and “low” groups based on pre-experiment characteristics that capture scale, experience, or market position. For sellers, the classification is based on size and tenure on the platform. For buyers, we use measures of online shopping experience and purchasing intensity. For products, we rely on indicators of category concentration, sales volume, and relative price. These groupings allow us to assess whether GenAI adoption disproportionately benefits smaller or less experienced sellers, less sophisticated buyers, and lower-volume or lower-priced products, thereby revealing how the technology shapes outcomes across different segments of the marketplace. Note that, depending on the context and data availability, some heterogeneity analyses cannot be implemented for certain workflows. We find consistent results across seller and buyer characteristics: disadvantaged groups, such as small sellers and inexperienced buyers, experience greater gains from GenAI-powered workflows. In contrast, the effects across product groups are more context-dependent.
@@ 5.1 Heterogeneous Effects Across Sellers
The platform hosts a highly diverse population of sellers, varying substantially in firm characteristics. We focus on sellers’ size and sophistication to investigate whether the productivity potential of GenAI vary between big versus small sellers. We classify sellers into high and low groups based on three pre-experiment metrics: annual sales value, years of operation on the platform (Operation Years), and the number of sub-accounts linked to the seller’s online store (# of Sub-Accounts). For each metric, sellers in the low group are generally small sellers, defined as meeting the following pre-experiment criteria: (i) accounting for the bottom 50% cumulative share of total sales when sellers are ranked by annual sales; (ii) having operated on the platform for fewer than five years; [^18] The platform closely monitors seller tenure, with five years of continuous operation serving as an important performance indicator. or (iii) maintaining fewer than three sub-accounts for their online store. [^19] According to the advice of the platform’s internal staff, small online stores are typically run by individuals or operate as mom-and-pop businesses, often with no more than two sub-accounts. Stores that have more sub-accounts generally employ additional staff to assist with tasks related to store operations, suggesting that these are more likely to be larger sellers. Table 7 presents the summarized results on seller heterogeneity, reporting only the percent change and significance levels for brevity. Additional results and details can be found in Appendix D.
@@
2 Marketing Sales 1.9% 2.1% 6.2% 0.7% 7.5% -0.8%
Push Message Conversion 3.3% 3.2%* 1.9% 3.8%** -0.1% 5.3%**
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
Table 7 shows that small sellers experience significant gains in both sales and conversion rates (Columns 2, 4, and 6), whereas effects for larger sellers tend to be smaller and statistically insignificant (Columns 1, 3, and 5). GenAI enhances the search algorithm’s ability to translate queries by understanding consumer intent and refining queries to improve semantic accuracy and clarity. This refinement allows the search engine to retrieve products more relevant to consumer needs, thereby reducing search frictions, improving matching efficiency, and ultimately increasing purchases for small sellers. These findings are consistent with prior work showing that lowering search costs or improving search quality fosters long-tail dynamics and expands market share for smaller and niche sellers [14, 5, 65, 69].
@@
A similar pattern, but only with respect to conversion rates, is observed in the platform’s advertising function, where only small sellers show significant gains. A likely explanation is that human drafters, constrained by time and attention, tend to reference or replicate the product characteristics of top sellers when composing marketing content. By contrast, GenAI enables the large-scale generation of customized marketing messages across a broad product portfolio, enhancing content differentiation for smaller sellers’ offerings and stimulating consumer purchases. These findings align with [60], who show that personalized recommendation systems disproportionately benefit smaller and less established sellers.
@@
We cannot assess seller heterogeneity in the Pre-sale Service Chatbot and Product Description experiments, as both were conducted exclusively on platform self-sold products with only a handful of platform-operated sellers, leaving insufficient variation for meaningful analysis. In the case of Product Description, as well as for experiments where granular data are unavailable (e.g., Chargeback Defense), the platform’s internal data science team provides descriptive evidence from third-party sellers to offer anecdotal insights into potential heterogeneity from GenAI. Specifically, third-party sellers are divided into quartiles based on pre-experiment annual sales. The share of products with no or limited descriptions is 1.2 times higher among bottom-quartile (small) sellers relative to top-quartile (large) sellers, while the share of unaddressed chargeback disputes is 1.7 times higher. These figures suggest that small sellers are particularly constrained by shortages of talent and resources, lacking the capacity to perform or outsource functions such as customer service, product content generation, or compliance management. Taken together, these patterns provide anecdotal support for the view that small sellers stand to benefit disproportionately from GenAI.
Overall, our findings indicate that the productivity gains from GenAI are disproportionately larger for small sellers, characterized by lower transaction volumes, shorter operational histories, and fewer sub-accounts. This pattern is consistent with prior research on GenAI adoption by individual workers, which shows that low-skilled and less experienced labor—often disadvantaged by earlier technological advances—derive greater benefits from GenAI [15, 43, 21, 54, 26, 22, 56]. Our results extend this insight to the domain of entrepreneurship on online retail platforms.

### v2

@@ 3.4 Data and Estimation
For the experiments operated at the consumer level, we recorded each consumer’s treatment status. We gathered consumers’ complete set of activities, including the number of product views (referred to as View), product clicks (Click), product orders (Order), and total expenditures on those orders (Sales). [^13] In the Search Query Refinement experiment, product views represent the number of products a consumer browses on the search results page, which displays a summarized collection of products immediately after a query search. In the Google Advertising Title experiment, product views refer to the number of views of advertised products within the Google Shopping tab. In the Search Query Refinement and Product Descriptions experiments, product clicks capture the number of times consumers clicked into product detail pages. In the Marketing Push Message experiment, product clicks reflect consumer clicks on push notifications, while in the Google Advertising Title experiment, they indicate clicks on advertised products. For comparability across workflows, our benchmark analysis focuses on two primary outcome measures: (i) total consumer expenditure (Sales), which serves as a revenue-based measure of retail productivity; and (ii) the conversion rate, a binary indicator for whether a consumer made a purchase during the experimental period, representing a proxy for changes in consumer experience. Both sales and conversion rate are widely used industry metrics. For the product-level experiment, we collected analogous data at the product level. Table 3 presents summary statistics for the key variables in each of the five experiments where granular data are available.
@@
Conversion Rate 0.04 0.21 0 0 1
Sales 0.51 4.56 0 0 2,942
@@
Click 0.017 0.131 0 0 1
Order 0.0018 0.053 0 0 6
Conversion Rate 0.0016 0.040 0 0 1
Sales 0.025 1.510 0 0 501
@@ 6 Chargeback Defense 15% defense success rate increase
The remaining four workflows with detailed data reported in Table 4 exhibit smaller effects on sales, ranging from a negative and statistically insignificant change to gains of up to 3%. Specifically, the Search Query Refinement application increased sales by 2.93% ($p<0.05$), while the Product Description generated a 2.05% ($p<0.05$) gain—still substantial effects for a platform of this scale and maturity. The Marketing Push Message workflow shows a positive yet not statistically significant improvement in sales (1.6%). This likely reflects the combination of a very low baseline conversion rate (only 0.14% of consumers make a purchase) and high variance in expenditures among converters, suggesting that broader implementation could provide sufficient power to detect treatment effects. [^15] Note that this case exhibits significant conversion rate increase (3.1%) in Table 4 as well as significant order increase (2.8%) in Appendix Table C5. It is also important to note that only a subset of the treatment group was exposed to GenAI-generated messages in this experiment, which may attenuate the observed treatment effects under partial exposure. By contrast, the Google Advertising Title workflow exhibits an insignificant negative effect. This pattern is consistent with the lack of fine-tuning of the GenAI model to the advertising context within the e-commerce domain, which led it to omit commonly used commercial keywords. An alternative explanation is a divergent delivery bias in ad distribution: Google’s advertising algorithm may have recognized and deprioritized AI-generated titles, reducing their visibility. Together, these findings underscore the importance of domain-specific fine-tuning or retraining of foundation models when applying GenAI to industry-specific tasks requiring specialized knowledge [27].
@@
These results provide new evidence on the potential of GenAI to enhance revenue-based productivity in online retail, thereby contributing to the broader debate on the economic consequences of GenAI adoption [56, 1]. A key distinction of our study is its focus on revenue outcomes rather than input-side efficiency gains, which makes our estimates not directly comparable to those from studies emphasizing labor productivity or input-side task efficiency. Among the seven deployments we examine, most deliver measurable performance gains, showing that GenAI can generate substantial improvements in firm outcomes under real-world operating conditions. Overall, the evidence points to a positive but heterogeneous impact of GenAI along the customer journey.
@@ 4.2 Mechanism
Most of the existing literature attributes the productivity-enhancing potential of GenAI to supply-side mechanisms, such as labor savings or efficiency improvements (e.g., reductions in the time required to complete a task). Our context and findings highlight an additional channel: GenAI can unlock productivity gains by enriching the demand-side consumer experience. We measure this mechanism using conversion rates (i.e., the likelihood that consumers complete a purchase at the extensive margin), which serve as a widely recognized industry standard for consumer satisfaction in the e-commerce context. Across the various workflows, we document significant increases in conversion rates ranging from 1% to 22% (Column 4 of Table 4), which in turn translate into higher output as measured by sales. Notably, we do not find any significant effects along the intensive margin. Table 5 shows that the average cart value among consumers who made at least one purchase (or among products purchased at least once) remains unchanged following GenAI adoption. In our setting, GenAI primarily drove market expansion by converting a larger share of consumers, rather than inducing existing buyers to purchase higher-priced or larger quantities of products.
@@
    The table presents the treatment effects of GenAI adoption on “Cart Value”. For workflows 1-4, “Cart Value” refers to the expenditure per consumer, conditional on the consumer making a purchase. For workflow 5, “Cart Value” refers to the sales per product, conditional on the product being purchased.
@@
    Column (1) reports the estimated coefficient, with standard errors in brackets. Column (2) reports % Change, which is calculated dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
This evidence indicates that the observed productivity gains arise from GenAI’s capacity to reduce market frictions—both by enabling new services and by improving existing ones—thereby enhancing the shopping experience and influencing purchasing behavior. For example, GenAI mitigates information asymmetries by providing relevant and timely assistance through pre-sale chatbots (a 21.7% increase in conversion rate, $p<0.01$) and by generating comprehensive and structured product descriptions (a 1.27% increase, $p<0.01$). It reduces search frictions and improves match quality by enhancing the translation and semantic comprehension of consumer queries (a 1.15% increase, $p<0.05$). In addition, GenAI enables personalization of marketing content by enabling large-scale generation of customized marketing messages across a broad product portfolio (a 3% increase, $p<0.05$). Finally, evidence from Chargeback Defense and Live Chat Translation indicates improvements consistent with enhanced experiences for sellers and consumers, respectively, including higher dispute success rates and greater customer satisfaction.
@@ 4.3 Aggregate Productivity Gains Across Workflows
1.64
1.31
2 Search Query Refinement 0.0648 40.6 2.63
3 Product Description 0.0104 52.1 0.54
4 Marketing Push Message 0.000402 365.0 0.15
Total (linear additivity) 4.96 (upper-bound)
@@
    Column (1) reports the absolute lifts in sales (treatment effects of GenAI) for each workflow from Table 4. Column (2) shows the factor used to annualize the workflow-specific estimates. Column (3) reports the annualized values for each workflow.
@@
Table 6 summarizes the key variables used in our calculations. Column 1 reports the AI-driven incremental value per consumer for each experimental workflow, reflecting the estimated average treatment effects (the absolute sales lift per consumer) reported in Column 1 of Table 4. Because each experiment had a different duration, Column 2 presents the time multiplier used to extrapolate these effects to an annual horizon. [^18] We assume that each experiment’s duration reflects the typical interval between treatment opportunities for a representative consumer on the platform. For example, a representative consumer is assumed to interact with the Pre-sale Service Chatbot about once every two months, whereas they could receive a Marketing Push Message daily. For instance, the Pre-sale Service Chatbot experiment spans two months, yielding a time multiplier of six (i.e., $12/2$). For each workflow, we obtain the annualized AI-driven incremental value per consumer in Column 3 by multiplying the estimated effect in Column 1 by the corresponding time multiplier in Column 2. This calculation assumes that the treatment effects observed during the experimental period remain constant over time, abstracting from potential amplification (e.g., greater engagement and purchases as GenAI-generated content enhances consumer satisfaction and platform loyalty) or attenuation (e.g., consumer dissatisfaction and product returns due to potential mismatches between AI-generated content and actual product characteristics).
@@
Finally, we aggregate the annualized estimates across the four workflows to obtain the total annual incremental value per consumer attributable to GenAI (see “Total” in Column 3). This aggregation assumes that effects across workflows are linearly additive and abstracts from potential cross-workflow interactions, such as cannibalization across touchpoints or expansion through synergies. Based on the four GenAI applications with positive sales effects, we estimate an annual incremental value of up to \$4.96 per consumer, which decreases to \$4.63 when using the lower-bound estimate from the Pre-sale Service Chatbot experiment. These effects represent roughly 5.5-6% of the per-user revenue growth observed in global e-commerce between 2023 and 2024 [59], highlighting the economic significance of these gains relative to broader industry trends.
@@
It is worth noting that these estimates capture only a partial and time-dependent view of the firm’s efforts to scale up GenAI across workflows. While in 2023 the platform applied GenAI to only a handful of workflows, by 2024 it was deployed in more than 40 applications and by 2025 in over 60. This rapid expansion is also reflected in the growth of API calls to the platform’s GenAI tools: in mid-2024, AI-related API requests averaged over 50 million per day, rising to more than 1 billion per day by mid-2025—a twentyfold increase. Hence, our point estimates should be interpreted with caution and in light of these rapid adoption trends, which nonetheless indicate that the firm anticipates substantial value from GenAI.
@@ 5 Heterogeneous Treatment Effects
To further shed light on the mechanism behind our results, we examine heterogeneity in treatment effects across sellers, buyers, and products. If GenAI primarily reduces frictions on both the demand and supply sides, one would expect relatively larger gains among participants with lower baseline capabilities—namely, smaller and newer sellers, less experienced buyers with limited platform engagement, and products in the long tail of sales or in less concentrated categories. We classify each dimension into “high” and “low” groups based on pre-experiment characteristics that capture scale, experience, or market position. For sellers, the classification is based on size and tenure on the platform. For buyers, we use measures of online shopping experience and purchasing intensity. For products, we rely on indicators of category concentration, sales volume, and relative price. These groupings allow us to assess whether GenAI adoption disproportionately benefits smaller or newer sellers, less sophisticated buyers, and products in the long tail or in less concentrated categories, thereby revealing how the technology shapes outcomes across different segments of the marketplace. Note that, depending on the context and data availability, some heterogeneity analyses cannot be implemented for certain workflows. We find consistent results across seller and buyer characteristics: disadvantaged groups, such as small sellers and inexperienced buyers, experience greater gains from GenAI-powered workflows. In contrast, the effects across product groups are more context-dependent.
@@ 5.1 Heterogeneous Effects Across Sellers
The platform hosts a highly diverse population of sellers, varying substantially in firm characteristics. We focus on sellers’ size and sophistication to investigate whether the productivity potential of GenAI vary between big versus small sellers. We classify sellers into high and low groups based on three pre-experiment metrics: annual sales value (Annual Sales), years of operation on the platform (Operation Years), and the number of sub-accounts linked to the seller’s online store (# of Sub-Accounts). For each metric, sellers in the low group are generally small sellers, defined as meeting the following pre-experiment criteria: (i) accounting for the bottom 50% cumulative share of total sales when sellers are ranked by annual sales; (ii) having operated on the platform for fewer than five years; [^19] The platform closely monitors seller tenure, with five years of continuous operation serving as an important performance indicator. or (iii) maintaining fewer than three sub-accounts for their online store. [^20] According to the advice of the platform’s internal staff, small online stores are typically run by individuals or operate as mom-and-pop businesses, often with no more than two sub-accounts. Stores that have more sub-accounts generally employ additional staff to assist with tasks related to store operations, suggesting that these are more likely to be larger sellers. Table 7 presents the summarized results on seller heterogeneity, reporting only the percent change and significance levels for brevity. Additional results and details can be found in Appendix D.
@@
2 Marketing Sales 1.8% 1.4% 5.1% 0.6% 6.8% -1.0%
Push Message Conversion 3.3% 2.7% 0.9% 3.8%** -0.1% 4.8%**
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
Table 7 shows that small sellers experience significant gains in both sales and conversion rates (Columns 2, 4, and 6), whereas effects for large sellers tend to be smaller and statistically insignificant (Columns 1, 3, and 5). GenAI enhances the search algorithm’s ability to translate queries by understanding consumer intent and refining queries to improve semantic accuracy and clarity. This refinement allows the search engine to retrieve products more relevant to consumer needs, thereby reducing search frictions, improving matching efficiency, and ultimately increasing purchases for small sellers. These findings are consistent with prior work showing that lowering search costs or improving search quality fosters long-tail dynamics and expands market share for smaller and niche sellers [14, 5, 65, 69].
@@
A similar pattern, but only with respect to conversion rates, is observed in the platform’s advertising function, where only small sellers show significant gains. A likely explanation is that human drafters, constrained by time and attention, tend to reference or replicate the product characteristics of top sellers when composing marketing content. By contrast, GenAI enables the large-scale generation of customized marketing messages across a broad product portfolio, enhancing content differentiation for small sellers’ offerings and stimulating consumer purchases. These findings align with [60], who show that personalized recommendation systems disproportionately benefit smaller and less established sellers.
@@
We cannot assess seller heterogeneity in the Pre-sale Service Chatbot and Product Description experiments, as both were conducted exclusively on platform self-sold products with only a handful of platform-operated sellers, leaving insufficient variation for meaningful analysis. In the case of Product Description, as well as for experiments where granular data are unavailable (e.g., Chargeback Defense), the platform’s internal data science team provides descriptive evidence from third-party sellers to offer anecdotal insights into potential heterogeneity in sellers’ responses to GenAI. Specifically, third-party sellers are divided into quartiles based on pre-experimental annual sales. The share of products with no or limited descriptions is 1.2 times higher among bottom-quartile (small) sellers relative to top-quartile (large) sellers, while the share of unaddressed chargeback disputes is 1.7 times higher. These figures suggest that small sellers are particularly constrained by shortages of talent and resources, lacking the capacity to perform or outsource functions such as customer service, product content generation, or compliance management. Taken together, these patterns provide anecdotal support for the view that smaller sellers stand to gain disproportionately more from GenAI.
Overall, our findings indicate that the productivity gains from GenAI are disproportionately larger for small sellers, characterized by lower transaction value, shorter operational histories, and fewer sub-accounts. This pattern is consistent with prior research on GenAI adoption by individual workers, which shows that low-skilled and less experienced labor—often disadvantaged by earlier technological advances—derive greater benefits from GenAI [15, 43, 21, 54, 26, 22, 56]. Our results extend this insight to the domain of entrepreneurship on online retail platforms.